Main INDEX
Monthly INDEX
PREV
NEXT
User name NIlanga
Log entry time 01:49:01 on May 2,2003
Entry number 99394
keyword=No target/12C target rates.
Last night when we started optics runs with the septum set at 225 A, the scaler rate for no target was about 4 kHz compared to with the 12C target about 44 kHz. (runs 1422 and 1423). In this case when I look at y_rot vs. ph_rot histogram under the same cut in x vs. y, for the 12C target case I see a very strong line which looks like the 12C target. In the no target case only a very faint shadow of this peak present.
But since then for every septum current setting we have both target positions, the rates for the two targets have been about the same. And the spectra look similar (no target setting seems to have the same strong line with some added junk) (see runs 1454 and 1455) plots attached to the next entry
FIGURE 1
FIGURE 2